Solve for x: logbasex(3x+10)=logbase7 49
Answer by edjones(8007)

log[x](3x+10)=log[7](49)
log[x](3x+10)=2
x^log[x](3x+10)=x^2
3x+10=x^2
x^2-3x-10=0
(x-5)(x+2)=0
x=5
.
Ed